# Few-Shot Adversarial Prompt Learning on Vision-Language Models

### v1

@@ Abstract
The vulnerability of deep neural networks to imperceptible adversarial perturbations has attracted widespread attention. Inspired by the success of vision-language foundation models, previous efforts achieved zero-shot adversarial robustness by aligning adversarial visual features with text supervision. However, in practice, they are still unsatisfactory due to several issues, including heavy adaptation cost, suboptimal text supervision, and uncontrolled natural generalization capacity. In this paper, to address these issues, we propose a few-shot adversarial prompt framework where adapting input sequences with limited data makes significant adversarial robustness improvement. Specifically, we achieve this by providing adversarially correlated text supervision that is end-to-end learned from adversarial examples. We also propose a novel training objective that enhances the consistency of multi-modal features while encourages differentiated uni-modal features between natural and adversarial examples. The proposed framework gives access to learn adversarial text supervision, which provides superior cross-modal adversarial alignment and matches state-of-the-art zero-shot adversarial robustness with only 1% training data.
[…]
@@ 1 Introduction
The seminal works (Szegedy et al., 2013; Goodfellow et al., 2014) reveal that adversarial examples (Goodfellow et al., 2014), consisting of malicious perturbations imperceptible to humans, can easily mislead state-of-the-art deep neural networks (DNNs) (Krizhevsky et al., 2012; He et al., 2016; Mahmood et al., 2021; Dong et al., 2023) into making incorrect predictions. This vulnerability limits the application of DNNs in safety-critical areas, such as medicine (Buch et al., 2018), healthcare (Finlayson et al., 2019), and autonomous driving (Tuncali et al., 2018).
@@
Human cognition is immune to the distribution variations induced by adversarial attacks, reflecting a fundamental difference between human and machine cognitive understanding. Humans primarily rely on semantic information (Zhang et al., 2021) from the context, while machines depend more on statistical distributional associations. Consequently, recent work (Mao et al., 2022) introduces text supervision in adversarial adaptation through foundational vision language models (VLMs) (Radford et al., 2021; Jia et al., 2021; Kim et al., 2021; Yao et al., 2021; Yuan et al., 2021; Li et al., 2022a; Yu et al., 2022b; Li et al., 2023a), enhancing adversarial robustness with improved semantic understanding. Specifically, they adapt visual prompts by aligning adversarial visual features with static text supervision from the CLIP model (Radford et al., 2021). By narrowing the gap in the probability distribution between adversarial text-image logits and the ground-truth label, they achieve zero-shot adversarial robustness in downstream tasks.
@@
However, although some progress has been made with the previous method, there are still three limitations to overcome before leveraging context to mitigate adversarial vulnerabilities. First, zero-shot adversarial robustness in downstream tasks stems from aligning image and text embeddings on large-scale generic datasets like the entire ImageNet (Deng et al., 2009) through adversarial adaptation, which necessitates a huge amount of time and computational resources. Second, static hand-crafted text prompts lack adversary-related hints, providing only content-related information while disregarding adversarial components. Finally, the current adaptation method only considers adversarial inputs while disregarding natural inputs. On the one hand, it fails to account for the relationship and distinctions between natural and adversarial examples, potentially leading to catastrophic forgetting of natural generalization during adversarial adaptation. Worse still, if there are distributional discrepancies in the downstream datasets, the constrained natural generalization could hinder the learning of robustness.
@@
To address these issues, we propose a Few-shot Adversarial Prompt learning (FAP) framework where pre-trained VLMs are adversarially adapted in a few-shot manner (Wang et al., 2021; Dong et al., 2022) with prompt learning (Lester et al., 2021; Liu et al., 2023b, a; Jia et al., 2022; Bahng et al., 2022; Zhou et al., 2022b; Zhang et al., 2023a; Li et al., 2024). This adapts the inputs rather than the parameters of the model. To the best of our knowledge, this is the first time to learn adversarial robustness from the perspective of few-shot prompt tuning. Due to the scarcity of data for establishing robust decision boundaries, the robust representations learned by existing adversarial visual prompt methods (Mao et al., 2022) are far from satisfactory. This leads us to rethink how to provide appropriate prompts for adversarial examples. Instead of using static hand-crafted text prompts, we propose to learn adversarially correlated text supervision end-to-end from adversarial examples. Moreover, we design a novel training objective that harmonizes the connection and distinction of natural and adversarial features from information across different modalities. That is, we force the multi-modal features of natural and adversarial inputs to be consistent while encouraging the differentiation between uni-modal embeddings.
@@
Before delving into details, we clearly summarize our contributions as follows:
We focus on a realistic and important research problem and discuss three major issues in previous adversarial prompt learning paradigms, potentially inspiring further improvements in this area.
To tackle these issues, we propose a novel adversarial few-shot prompt learning framework with learnable adversarial text supervision and an adversarial-aware prompt learning objective. This method is lightweight yet makes significant adversarial generalization.
We justify our claims through a series of experiments on 11 benchmark datasets covering multiple recognition tasks. The proposed method significantly outperforms state-of-the-art adversarial prompt learning methods in adversarial few-shot learning, adversarial zero-shot transfer, and adversarial base-to-new generalization settings. Comprehensive ablation studies and discussions are also provided.
@@ 2 Preliminary
CLIP recap. A pre-trained CLIP model typically includes an image encoder with learned parameters and a text encoder with learned parameters . Here we consider a -class classification problem for an image and its corresponding label . To perform zero-shot evaluation, is first divided into patches and converted into the patch embeddings . A class token is then appended to the patch sequence as . Afterward, the image encoder processes this embedded patch sequence with ViT (Dosovitskiy et al., 2020) blocks to produce the latent image feature representation . For the text branch, we prepare hand-craft prompts by appending the class name to a word template, such as ‘a photo of a {class}’. Subsequently, is tokenized and embedded as , where corresponds the -th class. The text encoder then encodes these work embeddings into the latent text feature representation . For zero-shot classification, the probability of the image in the -th class is
@@
where corresponds the text-image logits. We suggest readers check Zhou et al. (2022b) for more details about CLIP-based prompt learning.
Adversarial visual prompt. Adversarial prompt learning optimizes prompt tokens through adversarial training, enhancing model robustness in a relatively small adaptation cost without altering the pre-trained model. Mao et al. (2022) achieves this by adjusting the visual prompt of adversarial images in joint text-image feature space. Notably, owing to the application of text-image contrastive loss during the generation of adversarial examples, the adapted model reveals zero-shot adversarial robustness on downstream tasks. Formally, let be the input feature space with the infinity distance metric, where . Adversarial data falls in to close ball of radius centered at . That is, . The learnable image prompt is inserted to the visual patch embedding of , as . Then, adversarial data is generated by maximizing the text-image contrastive loss as
@@
Here, is defined as a text-image contrastive adversarial training (TeCoA) loss by Mao et al. (2022) that highlights adversarial text-image alignment.
@@ 3 Method
Overview. To address the limitations of previous methods, we propose FAP, a few-shot adversarial prompt learning framework. Our framework uses lightweight learnable prompts on the top of the pre-trained CLIP in a few-shot manner, as the case in natural prompt tuning (Zhou et al., 2022b). In more detail, we introduce learnable prompt tokens for adversarial examples, which allows the model to provide more appropriate text supervision that helps balance natural and adversarial generalization. Based on CLIP’s dual-encoder architecture, we further provide a novel training objective that guides the discrimination of natural and adversarial embeddings in uni-modal feature space. This promotes uni-modal divergence to incorporate an adversarial-aware mechanism, facilitating the learning of adversarial text supervision. The overview of the proposed framework is provided in Figure 1. Below, we discuss the FAP framework step by step.
@@ 3.1 Learnable Text Supervision for Adversarial Examples
When adapting the CLIP model, a slight change in wording could have a huge impact on performance (Zhou et al., 2022b). With the existence of adversarial examples, the situation has become worse. The distribution differences between natural and adversarial examples necessitate the design of specialized text supervision specifically for adversarial samples. Therefore, we introduce text prompt tokens that are end-to-end learned through adversarial examples.
Formally, our adversarial prompt learning is implemented on a few-shot subset , created by sampling examples from each of the classes in the original dataset. Learnable prompts consist of both visual and text branches, denoted as . The visual prompt token is incorporated into the image embedding, as observed in an adversarial visual prompt, while text prompt token is inserted into word embedding, as is the case in natural prompt learning. To preserve mutual synergy between visual and text branchs, is obtained from through linear projection , which can be denoted as .
Discussion. The proposed framework can be categorized as a cross-modal prompt, drawing inspiration from natural multi-modal prompt learning framework (Khattak et al., 2023a). However, we identify objective differences under adversarial task settings and make minimal adjustments for adversarial robustness. Specifically, Khattak et al. (2023a) employs deep prompt interaction in a text-to-image context, as hand-crafted prompts already provide suitable initial text supervision for clean examples. Additionally, the text-to-image projection acts as a dimensionality-increasing projection, preventing information loss. In contrast, the learning of text prompts in our adversarial setting relies on features from adversarial examples derived from the visual branch. Therefore, we adapt to adversarial prompt learning by simply reversing the projection direction to image-to-text interaction. We offer a comprehensive analysis of deep interaction settings in Appendix D.1.
@@ 3.2 Balancing Natural and Adversarial Generalization in Few-Shot Adversarial Prompt
For adapting the CLIP model to adversarial robustness tasks, the existing method (Mao et al., 2022) proposes the adversarial text-image contrastive (TeCoA) loss (4). This method minimizes the discrepancy between the distribution of adversarial text-image similarity and one-hot ground-truth labels. While this strategy effectively aligns text representations during adversarial adaptation, it potentially compromises the model’s generalization ability in specific recognition tasks under few-shot conditions.
The effectiveness of this method is contingent on the similarity between the downstream task’s distribution and the pre-trained representations. Specifically, if the downstream task aligns closely with the pre-trained representation, the CLIP model demonstrates favorable natural generalization. In such scenarios, employing additional learnable prompts for robustness adaptation proves beneficial. Conversely, a significant disparity between the distribution of the downstream task and the pre-trained representation poses a challenge. In such cases, the CLIP model inherently lacks natural generalization. Consequently, expecting the prompt tokens to concurrently learn both natural and robust generalization from a limited set of adversarial examples becomes an overly demanding task.
Balancing natural and adversarial generalization. Inspired by the success of TRADES (Zhang et al., 2019) in standard adversarial training, we propose a surrogate adversarial text-image contrastive loss that decouples the adversarial text-image contrastive loss into natural and adversarial terms. By encoding image and text embeddings with their respective transformer encoder and calculating similarity across modality, we have the natural and adversarial text-image logits: and , where . The learning objective can be stated as:
where denotes the Kullback–Leibler (KL) divergence and is a weight parameter. In Eq. (5), the first term encourages minimizing the natural error between the natural text-image similarity and label. The second term minimizes the boundary error by narrowing the distribution gap between natural and adversarial text-image similarity to ensure cross-modal adversarial consistency. We argue that a balanced two-term objective is crucial for downstream generalization, as this design alleviates the potential failure in robustness caused by discrepancies in natural generalization.
@@ 3.4 Overall Learning Objective
For better understanding, we describe our adversarial prompt learning and adversarial prompt testing pipeline in Appendix A. Additionally, we demonstrate the significant robustness gains our learning objective brings to other prompt designs through a case study in Appendix D.2. Different training objective designs and their experimental results can be found in Appendix D.3.
@@ 4.1 Setups
Baselines. To demonstrate the expertise of the proposed method, we employ the adversarial version of multiple commonly used prompt learning designs for comparison. Specifically, we compare our FAP with the following baselines that can be divided into two groups:
(1) Methods with hand-craft text supervision, which include zero-shot CLIP (Radford et al., 2021) and adversarial visual prompt method (AdvVP) (Mao et al., 2022).
(2) Methods with learnable text supervision, such as adversarial vision language prompts (AdvVLP) and adversarial multi-modal prompts (AdvMaPLe) (Khattak et al., 2023a). Note that we primarily focus on methods that add a learnable prompt on top of the AdvVP scheme, and the effects of pure text prompts (Zhou et al., 2022b) under the adversarial setting (AdvTP) are attached in AppendixD.9.
The technical details about these baselines are provided in Appendix C.1, and the static prompt templates for each dataset can be found in Appendix C.2.
Datasets. To evaluate the proposed method, we align with previous works (Zhou et al., 2022b, a) and utilize 11 diverse image recognition datasets that span multiple vision tasks. Specifically, the datasets include two generic object datasets: ImageNet (Deng et al., 2009) and Caltech101 (Fei-Fei et al., 2004); a texture recognition dataset: DTD (Cimpoi et al., 2014); five fine-grained object recognition datasets: FGVCAircraft (Maji et al., 2013), OxfordPets (Parkhi et al., 2012), Flowers102 (Nilsback & Zisserman, 2008), Food101 (Bossard et al., 2014), and StanfordCars (Krause et al., 2013); a scene recognition dataset: SUN397 (Xiao et al., 2010); an action recognition dataset: UCF101 (Soomro et al., 2012); and a satellite image classification dataset: EuroSAT (Helber et al., 2019).
Implementation details. We conduct experiments on the ViT-B/32 CLIP architecture and report the average results over three random seeds. All models are trained for 5 epochs in cross-dataset evaluation and 10 epochs for other benchmark settings by using an SGD optimizer with a momentum of 0.9. The initial learning rate is set at 0.0035. We apply a cosine learning rate scheduler and a warm-up strategy during the first epoch. For adversarial prompt learning, we use token prompts of size 2 in both the vision and text branches across the first 9 transformer blocks. Attacks are generated under threat model through a 2-step PGD attack, with a perturbation boundary and a step size , following the methodologies outlined in (Mao et al., 2022). The adversarial robustness is evaluated using a 100-step PGD attack. Note that we explore alternative CLIP architectures in Appendix D.4, different attack strengths in Appendix D.5, and various choices of adversarial robustness evaluation method in Appendix D.6.

### v2

@@ Abstract
The vulnerability of deep neural networks to imperceptible adversarial perturbations has attracted widespread attention. Inspired by the success of vision-language foundation models, previous efforts achieved zero-shot adversarial robustness by aligning adversarial visual features with text supervision. However, in practice, they are still unsatisfactory due to several issues, including heavy adaptation cost, suboptimal text supervision, and uncontrolled natural generalization capacity. In this paper, to address these issues, we propose a few-shot adversarial prompt framework where adapting input sequences with limited data makes significant adversarial robustness improvement. Specifically, we achieve this by providing adversarially correlated text supervision that is end-to-end learned from adversarial examples. We also propose a novel training objective that enhances the consistency of multi-modal features while encourages differentiated uni-modal features between natural and adversarial examples. The proposed framework gives access to learn adversarial text supervision, which provides superior cross-modal adversarial alignment and matches state-of-the-art zero-shot adversarial robustness with only 1% training data. Code is available at: https://github.com/lionel-w2/FAP.
@@ 1 Introduction
The seminal works [1, 2] reveal that adversarial examples [2], consisting of malicious perturbations imperceptible to humans, can easily mislead state-of-the-art deep neural networks (DNNs) [3, 4, 5, 6] into making incorrect predictions. This vulnerability limits the application of DNNs in safety-critical areas, such as medicine [7], healthcare [8], and autonomous driving [9].
@@
Human cognition is immune to the distribution variations induced by adversarial attacks, reflecting a fundamental difference between human and machine cognitive understanding. Humans primarily rely on semantic information [10] from the context, while machines depend more on statistical distributional associations. Consequently, recent work [11] introduces text supervision in adversarial adaptation through foundational vision language models (VLMs) [12, 13, 14, 15, 16, 17, 18, 19], enhancing adversarial robustness with improved semantic understanding. Specifically, they adapt visual prompts by aligning adversarial visual features with static text supervision from the CLIP model [12]. By narrowing the gap in the probability distribution between adversarial text-image logits and the ground-truth label, they achieve zero-shot adversarial robustness in downstream tasks.
@@
However, although some progress has been made with the previous method, there are still three limitations to overcome before leveraging context to mitigate adversarial vulnerabilities. First, zero-shot adversarial robustness in downstream tasks stems from aligning image and text embeddings on large-scale generic datasets like the entire ImageNet [20] through adversarial adaptation, which necessitates a huge amount of time and computational resources. Second, static hand-crafted text prompts lack adversary-related hints, providing only content-related information while disregarding adversarial components. Finally, the current adaptation method only considers adversarial inputs while disregarding natural inputs. On the one hand, it fails to account for the relationship and distinctions between natural and adversarial examples, potentially leading to catastrophic forgetting of natural generalization during adversarial adaptation. Worse still, if there are distributional discrepancies in the downstream datasets, the constrained natural generalization could hinder the learning of robustness.
@@
To address these issues, we propose a Few-shot Adversarial Prompt learning (FAP) framework where pre-trained VLMs are adversarially adapted in a few-shot manner [21, 22] with prompt learning [23, 24, 25, 26, 27, 28]. This adapts the inputs rather than the parameters of the model. To the best of our knowledge, this is the first time to learn adversarial robustness from the perspective of few-shot prompt tuning. Due to the scarcity of data for establishing robust decision boundaries, the robust representations learned by existing adversarial visual prompt methods [11] are far from satisfactory. This leads us to rethink how to provide appropriate prompts for adversarial examples. Instead of using static hand-crafted text prompts, we propose to learn adversarially correlated text supervision end-to-end from adversarial examples. Moreover, we design a novel training objective that harmonizes the connection and distinction of natural and adversarial features from information across different modalities. That is, we force the multi-modal features of natural and adversarial inputs to be consistent while encouraging the differentiation between uni-modal embeddings.
@@
Before delving into details, we clearly summarize our contributions as follows. (1) We focus on a realistic and important research problem and discuss three major issues in previous adversarial prompt learning paradigms, potentially inspiring further improvements in this area. (2) To tackle these issues, we propose a novel adversarial few-shot prompt learning framework with learnable adversarial text supervision and an adversarial-aware prompt learning objective. This method is lightweight yet makes significant adversarial generalizations. (3) We justify our claims through a series of experiments on 11 benchmark datasets covering multiple recognition tasks. The proposed method significantly outperforms state-of-the-art adversarial prompt learning methods in adversarial few-shot learning, adversarial zero-shot transfer, and adversarial base-to-new generalization settings. Comprehensive ablation studies and discussions are also provided in Section 4.3 and Appendix D.
@@ 2 Preliminary
CLIP recap. A pre-trained CLIP model typically includes an image encoder with learned parameters and a text encoder with learned parameters . Here we consider a -class classification problem for an image and its corresponding label . To perform zero-shot evaluation, is first divided into patches and converted into the patch embeddings . A class token is then appended to the patch sequence as . Afterward, the image encoder processes this embedded patch sequence with ViT [29] blocks to produce the latent image feature representation . For the text branch, we prepare hand-craft prompts by appending the class name to a word template, such as ‘a photo of a {class}’. Subsequently, is tokenized and embedded as , where corresponds the -th class. The text encoder then encodes these work embeddings into the latent text feature representation . For zero-shot classification, the probability of the image in the -th class is
@@
where corresponds the text-image logits. We suggest readers check Zhou et al. [28] for more details about CLIP-based prompt learning.
Adversarial visual prompt. Adversarial prompt learning optimizes prompt tokens through adversarial training, enhancing model robustness in a relatively small adaptation cost without altering the pre-trained model. Mao et al. [11] achieves this by adjusting the visual prompt of adversarial images in joint text-image feature space. Notably, owing to the application of text-image contrastive loss during the generation of adversarial examples, the adapted model reveals zero-shot adversarial robustness on downstream tasks. Formally, let be the input feature space with the infinity distance metric, where . Adversarial data falls in to close ball of radius centered at . That is, . The learnable image prompt is inserted to the visual patch embedding of , as . Then, adversarial data is generated by maximizing the text-image contrastive loss as
@@
Here, is defined as a text-image contrastive adversarial training (TeCoA) loss by Mao et al. [11] that highlights adversarial text-image alignment.
@@ 3 Method
Overview. To address the limitations of previous methods, we propose FAP, a few-shot adversarial prompt learning framework. Our framework uses lightweight learnable prompts on the top of the pre-trained CLIP in a few-shot manner, as the case in natural prompt tuning [28]. In more detail, we introduce learnable prompt tokens for adversarial examples, which allows the model to provide more appropriate text supervision that helps balance natural and adversarial generalization. Based on CLIP’s dual-encoder architecture, we further provide a novel training objective that guides the discrimination of natural and adversarial embeddings in uni-modal feature space. This promotes uni-modal divergence to incorporate an adversarial-aware mechanism, facilitating the learning of adversarial text supervision. The overview of the proposed framework is provided in Figure 1. Below, we discuss the FAP framework step by step.
@@ 3.1 Learnable Text Supervision for Adversarial Examples
When adapting the CLIP model, a slight change in wording could have a huge impact on performance [28]. With the existence of adversarial examples, the situation has become worse. The distribution differences between natural and adversarial examples necessitate the design of specialized text supervision specifically for adversarial samples. Therefore, we introduce text prompt tokens that are end-to-end learned through adversarial examples.
Formally, our adversarial prompt learning is implemented on a few-shot subset , created by sampling examples from each of the classes in the original dataset. Learnable prompts consist of both visual and text branches, denoted as . The visual prompt token is incorporated into the image embedding, as observed in an adversarial visual prompt, while text prompt token is inserted into word embedding, as is the case in natural prompt learning. To preserve mutual synergy between visual and text branchs, is obtained from through linear projection , which can be denoted as . The proposed framework can be categorized as a cross-modal prompt [30] with minimal modification for adversarial robustness tasks. We offer a comprehensive analysis of the prompt design in Section 4.3.
@@ 3.2 Balancing Natural and Adversarial Generalization in Few-Shot Adversarial Prompt
For adapting the CLIP model to adversarial robustness tasks, the existing method [11] proposes the TeCoA loss (Eq.(4)). This method minimizes the discrepancy between the distribution of adversarial text-image similarity and one-hot ground-truth labels. While this strategy effectively aligns text representations during adversarial adaptation, it potentially compromises the model’s generalization ability in specific recognition tasks under few-shot conditions.
The method’s effectiveness depends on the similarity between the downstream task’s distribution and the pre-trained representations. When the downstream task closely aligns with the pre-trained representation, the CLIP model shows preferable natural generalization, and adding learnable prompts for robustness adaptation is advantageous. However, a significant mismatch between the downstream distribution and pre-trained representations challenges the CLIP model’s natural generalization capabilities. In such cases, expecting prompt tokens to learn both natural and robust generalization from a few adversarial examples is overly ambitious.
Balancing natural and adversarial generalization. Inspired by the success of TRADES [31] in standard adversarial training, we propose a surrogate adversarial text-image contrastive loss that decouples the adversarial text-image contrastive loss into natural and adversarial terms. By encoding image and text embeddings with their respective transformer encoder and calculating similarity across modality, we have the natural and adversarial text-image logits: and , where . The learning objective can be stated as:
where denotes the Kullback–Leibler (KL) divergence and is a weight parameter. In Eq. (5), the first term encourages minimizing the natural error between the natural text-image similarity and label. The second term minimizes the boundary error by narrowing the distribution gap between natural and adversarial text-image similarity to ensure cross-modal adversarial consistency. Note that a balanced two-term objective is crucial for downstream generalization, as this design alleviates the potential failure in robustness caused by discrepancies in natural generalization. We provide more analysis on the natural generalization gap in Appendix D.2.
@@ 3.4 Overall Learning Objective
3.5 Intuition behind Objective Design
Our learning objective highlights the differentiated processing of features under different modalities, in which we introduce an additional adversarial-aware mechanism with uni-modal image features. We discuss the intuition behind the design concept. We visualize the uni-modal embedding to demonstrate the impact of the adversarial-aware mechanism on the model’s feature learning.
In Figure 2(a), we find that certain adversarial embeddings closely resemble natural examples. This suggests that the consistency of cross-modal features between natural and adversarial examples arises from the model’s tendency to minimize loss by generating minimal adversarial perturbations. These exceedingly small perturbations do not effectively promote robust learning. In contrast, the adversarial-aware mechanism clearly separates the natural and adversarial embeddings in Figure 2(b), preventing the minimal perturbation shortcut and guiding the model to recognize the differences between natural and adversarial image embeddings.
For better understanding, we discuss different training objective designs and their results in Section 4.3 and describe our adversarial prompt learning and adversarial prompt testing pipeline in Appendix A. Additionally, we demonstrate the significant robustness gains our learning objective brings to other prompt designs through a case study. More details can be checked in Appendix D.4.
@@ 4.1 Setups
Baselines. To demonstrate the expertise of the proposed method, we employ the adversarial version of multiple commonly used prompt learning designs for comparison. We categorize our baselines into two groups: (1) Methods using hand-crafted text supervision, such as zero-shot CLIP [12] and AdvVP [11]. (2) Methods utilizing learnable text prompts, including AdvVLP and AdvMaPLe [30]. Note that we primarily focus on learnable prompts extending the AdvVP framework. Details on pure text prompt effects in adversarial settings (AdvTP) [28] are discussed in Appendix D.11. Additional information about these methods and static prompt templates for each dataset are provided in Appendices C.1 and C.2, respectively.
Datasets. To evaluate the proposed method, we align with previous works [28, 33] and utilize 11 diverse image recognition datasets that span multiple vision tasks. Specifically, the datasets include two generic object datasets: ImageNet-1K [20] and Caltech101 [32]; a texture recognition dataset: DTD [34]; five fine-grained object recognition datasets: FGVCAircraft [35], OxfordPets [36], Flowers102 [37], Food101 [38], and StanfordCars [39]; a scene recognition dataset: SUN397 [40]; an action recognition dataset: UCF101 [41]; and a satellite image classification dataset: EuroSAT [42].
Implementation details. We conduct experiments on the ViT-B/32 CLIP architecture and report the average results over three random seeds. All models are trained for 5 epochs in cross-dataset evaluation and 10 epochs for other benchmark settings by using an SGD optimizer with a momentum of 0.9. The initial learning rate is set at 0.0035. We apply a cosine learning rate scheduler and a warm-up strategy during the first epoch. For adversarial prompt learning, we use token prompts of size 2 in both the vision and text branches across the first 9 transformer blocks. Attacks are generated under threat model through a 2-step PGD attack, with a perturbation boundary and a step size , following the methodologies outlined in [11]. The adversarial robustness is evaluated using a 100-step PGD attack.
Note that due to the limited space of the main paper, we provide comprehensive evaluations, including cross-dataset evaluation (Appendix D.1), the comparison with AdvMaPLe (Appendix D.3), alternative CLIP architectures (Appendix D.5), different attack strengths (Appendix D.6), various choices of adversarial robustness evaluation methods (Appendix D.7), and different training-time attack generation (Appendix D.8).